# Extremal metrics involving scalar curvature

### v2

@@ 6.1 Other rigidity results
Theorem 6.1 ([2, Main Theorem]).
@@
Similarly, from [2, Theorem 4], for any Lipschitz map $f:M\rightarrow\mathbb{S}^{n}$ with $\mathrm{deg}(f)\neq 0$,
@@ 6.2 For singular metrics
Next, we mention that the scalar minimum functional $R_{min}$ (Section 2.1) and a generalized definition of scalar curvature bounded below. Let $M^{n}$ be a smooth closed $n$-manifold and $\kappa\in\mathbb{R}$ a constant. Let $C^{0}_{met}(M,\kappa)$ denote the $C^{0}$-completion of $C^{2}$-metrics whose scalar curvature is bounded below by $\kappa$ in the conventional sense. That is, a $C^{0}$-metric $g$ is in $C^{0}_{met}(M,\kappa)$ if and only if there exists a sequences of $C^{2}$-metrics $g_{i}$ on $M$ such that $g_{i}$ converges uniformly to $g$ and satisfies $R(g_{i})\geq\kappa$. From the observation after Remark 1.10 of [5] and Theorem 1.7 of the same paper, one can observe that a $C^{0}$-metric $g$ is in $C^{0}_{met}(M,\kappa)$ if and only if
@@
Here, “$\limsup_{C^{2}\ni h\rightarrow g,C^{0}}R_{min}(h)\geq\kappa$” means that
@@
then this is equivalent to that for any $\varepsilon>0$ there is $\delta>0$ such that
@@
$Y(M)=\sup_{h\in\mathcal{M}}R_{min}(h)\cdot\mathrm{Vol}(M,g)^{2/n}.$
Assume that there is a sequence of $C^{2}$-metrics $g_{i}$ on $M$ such that $g_{i}\overset{C^{0}}{\longrightarrow}g$ and
@@
Therefore, $g\in C^{0}_{met}\left(M,Y(M)\cdot\mathrm{Vol}(M,g)^{-2/n}\right)$. Of course, a typical example of $(g_{i})$ that satisfies only (26) is a sequence of solutions to the Yamabe problem on each conformal class $[g_{i}]$.
Related to this observation, it is interesting to explore some relations between variational properties of $R_{min}$ in a topology weaker than $C^{2}$ and singular Yamabe metrics or other extremal metrics.
@@ 6.3 About our assumptions
Question 6.4.
In our theorems 1.3, 1.4, 1.5 and 1.6, can we weaken the assumption that “$\|\overset{\circ}{\mathrm{Ric}_{g}}\|_{g}(x)\neq 0$ for all $x\in M$” to that “the metric $g$ is not an Einstein metric”?
@@
Several examples of manifolds that admit no Einstein metric are known. On the other hand, every closed manifold of dimension $\geq 3$ admits a metric with constant negative scalar curvature. Therefore, such a manifold of dimension $\geq 3$ always admits a metric of non-Einstein negative constant scalar curvature. Moreover, Matsuo [24, Corollary 2] proved that there exists a non-Ricci-flat scalar-flat metric on every closed manifold of dimension $\geq 3$ with positive Yamabe invariant. However, one cannot distinguish whether or not the norm of the traceless Ricci tensor of each such metric has a positive lower bound on the whole manifold.
Question 6.6.
Are there any examples that satisfy the assumptions of Theorem 1.7 or 1.8 (especially the scalar non-flat case)?
@@ 7 Appendix
Here, ¯$\bar{\nabla}_{i}g_{jk}$ denotes the expression of $\nabla^{\bar{g}}g$ in terms of the local coordinates.
@@
¯$\begin{split}\bar{\nabla}_{i}g_{lj}&=\partial_{i}g_{lj}-g_{pj}\bar{\Gamma}^{p}_{il}-g_{lp}\bar{\Gamma}^{p}_{ij}\\ \bar{\nabla}_{j}g_{il}&=\partial_{j}g_{il}-g_{pl}\bar{\Gamma}^{p}_{ji}-g_{ip}\bar{\Gamma}^{p}_{jl}\\ -\bar{\nabla}_{l}g_{ij}&=-\partial_{l}g_{ij}+g_{pj}\bar{\Gamma}^{p}_{il}+g_{ip}\bar{\Gamma}^{p}_{lj}.\end{split}$
Taking the sum of these both sides, we get
@@
$\begin{split}W^{p}_{ij}W^{k}_{kp}&=(\Gamma^{p}_{ij}-\bar{\Gamma}^{p}_{ij})(\Gamma^{k}_{kp}-\bar{\Gamma}^{k}_{kp})=\Gamma^{p}_{ij}\Gamma^{k}_{kp}-\underset{(1)}{\underline{\bar{\Gamma}^{p}_{ij}\Gamma^{k}_{kp}}}-\underset{(2)}{\underline{\Gamma^{p}_{ij}\bar{\Gamma}^{k}_{kp}}}+\underset{(3)}{\underline{\bar{\Gamma}^{p}_{ij}\bar{\Gamma}^{k}_{kp}}}\\ -W^{p}_{kj}W^{k}_{ip}&=-(\Gamma^{p}_{kj}-\bar{\Gamma}^{p}_{kj})(\Gamma^{k}_{ip}-\bar{\Gamma}^{k}_{ip})=-\Gamma^{p}_{kj}\Gamma^{k}_{ip}+\underset{(4)}{\underline{\bar{\Gamma}^{p}_{kj}\Gamma^{k}_{ip}}}+\underset{(6)}{\underline{\Gamma^{p}_{kj}\bar{\Gamma}^{k}_{ip}}}-\underset{(5)}{\underline{\bar{\Gamma}^{p}_{kj}\bar{\Gamma}^{k}_{ip}}}\\ \bar{\nabla}_{k}W^{k}_{ij}&=\bar{\nabla}_{k}(\Gamma^{k}_{ij}-\bar{\Gamma}^{k}_{ij})\\ &=\partial_{k}\Gamma^{k}_{ij}-\partial_{k}\bar{\Gamma}^{k}_{ij}+\underset{(2)}{\underline{\Gamma^{p}_{ij}\bar{\Gamma}^{k}_{kp}}}-\underset{(7)}{\underline{\Gamma^{k}_{pj}\bar{\Gamma}^{p}_{ik}}}-\underset{(4)}{\underline{\Gamma^{k}_{ip}\bar{\Gamma}^{p}_{kj}}}-\bar{\Gamma}^{p}_{ij}\bar{\Gamma}^{k}_{kp}+\underset{(8)}{\underline{\bar{\Gamma}^{k}_{pj}\bar{\Gamma}^{p}_{ik}}}+\underset{(5)}{\underline{\bar{\Gamma}^{k}_{ip}\bar{\Gamma}^{p}_{kj}}}\\ -\bar{\nabla}_{i}W^{k}_{kj}&=-\bar{\nabla}_{i}(\Gamma^{k}_{kj}-\bar{\Gamma}^{k}_{kj})\\ &=-\partial_{i}\Gamma^{k}_{kj}+\partial_{i}\bar{\Gamma}^{k}_{kj}-\underset{(6)}{\underline{\Gamma^{p}_{kj}\bar{\Gamma}^{k}_{ip}}}+\underset{(7)}{\underline{\Gamma^{k}_{pj}\bar{\Gamma}^{p}_{ki}}}+\underset{(1)}{\underline{\Gamma^{k}_{kp}\bar{\Gamma}^{p}_{ij}}}+\bar{\Gamma}^{p}_{kj}\bar{\Gamma}^{k}_{ip}-\underset{(8)}{\underline{\bar{\Gamma}^{k}_{pj}\bar{\Gamma}^{p}_{ki}}}-\underset{(3)}{\underline{\bar{\Gamma}^{k}_{kp}\bar{\Gamma}^{p}_{ij}}}.\end{split}$
@@
¯$\begin{split}\bar{\nabla}_{k}W^{k}_{ij}&-\bar{\nabla}_{i}W^{k}_{kj}+W^{p}_{ij}W^{k}_{kp}-W^{p}_{kj}W^{k}_{ip}\\ &=\partial_{k}\Gamma^{k}_{ij}-\partial_{i}\Gamma^{k}_{kj}+\Gamma^{p}_{ij}\Gamma^{k}_{kp}-\Gamma^{p}_{kj}\Gamma^{k}_{ip}\\ &~{}~{}~{}~{}~{}-(\partial_{k}\bar{\Gamma}^{k}_{ij}-\partial_{i}\bar{\Gamma}^{k}_{kj}+\bar{\Gamma}^{p}_{ij}\bar{\Gamma}^{k}_{kp}-\bar{\Gamma}^{p}_{kj}\bar{\Gamma}^{k}_{ip})\\ &=R_{ij}-\bar{R}_{ij}.\end{split}$
@@
Since ¯$\bar{\nabla}\bar{g}=0$, we have
@@
where the term $g^{-1}*g^{-1}*g^{-1}*\bar{\nabla}h*\bar{\nabla}h$ is a contraction of three copies of $g^{-1}$ (i.e., $g$ with raised indices) and two copies of ¯$\bar{\nabla}h=\bar{\nabla}g$. And, the term $(\bar{g}+h)^{-1}h\bar{g}^{-1}h\bar{g}^{-1}*\mathrm{Ric}_{\bar{g}}$ is the trace of $\mathrm{Ric}_{\bar{g}}$ with respect to $\left((\bar{g}+h)^{-1}h\bar{g}^{-1}h\bar{g}^{-1}\right)^{-1}$. Note that $\bar{g}+h$ is positive definite if $\|h\|_{\bar{g}}$ is small enough. Indeed, this formula follows by taking the both sides of (27) in Proposition 7.2 with respect to $g=\bar{g}+h$ and using

### v3

@@ 6.1 Other rigidity results
Theorem 6.1 ([3, Main Theorem]).
@@
Similarly, from [3, Theorem 4], for any Lipschitz map $f:M\rightarrow\mathbb{S}^{n}$ with $\mathrm{deg}(f)\neq 0$,
@@
6.2 A sufficient condition for a metric to be a positive Yamabe metric
Theorem 6.2 ([14]).
Let $M$ be a closed manifold and $g$ a Yamabe metric in its conformal class with $R_{g}>0$. Assume that a metric $h$ on $M$ has a positive constant scalar curvature and satisfies
$R_{g}\cdot g\geq R_{h}\cdot h~~\mathrm{on}~M.$ (24)
Then, $h$ is also a Yamabe metric in its conformal class. Moreover, if the inequality in (24) is strict, then $h$ is a unique Yamabe metric in its conformal class.
This type of sufficient condition is also known for other types of Yamabe metrics ([12, 13]).
On the other hand, according to [1], if $g$ is a strongly stable unique non-negative Yamabe metric with unit volume in its conformal class (then every metric sufficiently $C^{\infty}$-close to $g$ also contains the unique Yamabe metric in its conformal class), then for any $(0,2)$-tensor $h$ with $\mathrm{tr}_{g}h=0$, we have
$R_{\gamma_{t}}\cdot\gamma_{t}-R_{g}\cdot g=R_{g}(\gamma_{t}-g)+t\left(-\int_{M}\langle\overset{\circ}{\mathrm{Ric}}_{g},h\rangle\,d\mathrm{vol}_{g}\right)\gamma_{t}+o(t)$
for all sufficiently small $t>0$. Here, $\gamma_{t}\in[g_{t}:=g+th]$ is the unique Yamabe metric in its conformal class with unit volume. Since $R_{g}=\mathrm{const}$, $h=-\overset{\circ}{\mathrm{Ric}}_{g}$ is a tt-tensor with respect to $g$. Moreover, since $\gamma_{t}$ is the unique Yamabe (hence constant scalar curvature) metric in $\gamma_{t}]$ for all sufficiently small $t>0$,
$\gamma_{t}-g=t\,\mathrm{pr}_{TT}(h)+o(t)=th+o(t),$
where $\mathrm{pr}_{TT}$ is the projection onto the subspace consisting of all tt-tensors in $T_{g}\mathcal{M}_{1}$. Here, $\mathcal{M}_{1}$ is the space of all Riemannain metric on $M$ with unit volume and $T_{g}\mathcal{M}_{1}$ is the tangent space of $\mathcal{M}_{1}$ at the metric $g$. Therefore, we have
$\begin{split}R_{\gamma_{t}}\cdot\gamma_{t}-R_{g}\cdot g&=t\left(\|\overset{\circ}{\mathrm{Ric}}_{g}\|^{2}_{L^{2}(M,g)}g-(R_{g}+O(t))\,\overset{\circ}{\mathrm{Ric}}_{g}\right)+o(t)\\ &=t\left(\left(\|\overset{\circ}{\mathrm{Ric}}_{g}\|^{2}_{L^{2}(M,g)}+\frac{R^{2}_{g}}{n}\right)g-R_{g}\cdot\mathrm{Ric}_{g}\right)+o(t).\end{split}$
From the above, we ask the following question.
Question 6.3.
Let $M$ be a smooth manifold of positive Yamabe invariant $Y(M)>0$. For a strongly stable non-Einstein unique Yamabe metric $h$ on $M$ with unit volume, is there a positive Yamabe metric $g$ with unit volume for which the inequality in (24) is strict?
By [25, Corollary 2], for every closed manifold $M$ of dimension $\geq 3$ with positive Yamabe invariant, there is a non-Ricci-flat scalar-flat metric $g$ on $M$. Thus, the above argument with $h=\mathrm{Ric}_{g}$ for such $g$ implies that $R_{\gamma_{t}}>0$ for any sufficiently small $t>0$ where $\gamma_{t}\in[g+t\,\mathrm{Ric}_{g}]$ is the unique unit-volume Yamabe metric in its conformal class. Therefore, in particular, such a metric $g$ is not a local maximizer of the functional : $\mathcal{M}_{1}\ni h\mapsto Y(M,[h])$.
@@ 6.3 For singular metrics
Next, we mention that the scalar minimum functional $R_{min}$ (Section 2.1) and a generalized definition of scalar curvature bounded below. Let $M^{n}$ be a smooth closed $n$-manifold and $\kappa\in\mathbb{R}$ a constant. Let $C^{0}_{met}(M,\kappa)$ denote the $C^{0}$-completion of $C^{2}$-metrics whose scalar curvature is bounded below by $\kappa$ in the conventional sense. That is, a $C^{0}$-metric $g$ is in $C^{0}_{met}(M,\kappa)$ if and only if there exists a sequence of $C^{2}$-metrics $g_{i}$ on $M$ such that $g_{i}$ converges uniformly to $g$ and satisfies $R(g_{i})\geq\kappa$. From the observation after Remark 1.10 of [6] and Theorem 1.7 of the same paper, one can observe that a $C^{0}$-metric $g$ is in $C^{0}_{met}(M,\kappa)$ if and only if
@@
Here, $\limsup_{C^{2}\ni h\rightarrow g,C^{0}}R_{min}(h)\geq\kappa$” means that
@@
then this is equivalent to the property that for any $\varepsilon>0$ there is $\delta>0$ such that
@@
$Y(M)=\sup_{h\in\mathcal{M}}R_{min}(h)\cdot\mathrm{Vol}(M,h)^{2/n}.$
Assume that there is a sequence of $C^{2}$-metrics $g_{i}$ on $M$ such that $g_{i}\overset{C^{0}}{\longrightarrow}g\in C^{0}$ and
@@
Therefore, $g\in C^{0}_{met}\left(M,Y(M)\cdot\mathrm{Vol}(M,g)^{-2/n}\right)$. Of course, a typical example of $(g_{i})$ that satisfies only (25) is a sequence of solutions to the Yamabe problem on each conformal class $g_{i}]$.
In relation to this observation, it is interesting to explore some relations between the variational properties of $R_{min}$ and singular Yamabe metrics or other extremal metrics in a topology weaker than $C^{2}$.
@@ 6.4 About our assumptions
In our theorems 1.3, 1.4, 1.5 and 1.6, can we weaken the assumption that $\|\overset{\circ}{\mathrm{Ric}_{g}}\|_{g}(x)\neq 0$ for all $x\in M$” to that “the metric $g$ is not an Einstein metric”?
Question 6.6.
@@
Several examples of manifolds that admit no Einstein metric are known. On the other hand, every closed manifold of dimension $\geq 3$ admits a metric with constant negative scalar curvature. Therefore, such a manifold of dimension $\geq 3$ always admits a metric of non-Einstein negative constant scalar curvature. Moreover, Matsuo [25, Corollary 2] proved that there exists a non-Ricci-flat scalar-flat metric on every closed manifold of dimension $\geq 3$ with positive Yamabe invariant. However, one cannot distinguish whether the norm of the traceless Ricci tensor of each such metric has a positive lower bound on the whole manifold.
@@ 7 Appendix
Here, $\bar{\nabla}_{i}g_{jk}$ denotes the expression of $\nabla^{\bar{g}}g$ in terms of the local coordinates.
@@
$\begin{split}\bar{\nabla}_{i}g_{lj}&=\partial_{i}g_{lj}-g_{pj}\bar{\Gamma}^{p}_{il}-g_{lp}\bar{\Gamma}^{p}_{ij}\\ \bar{\nabla}_{j}g_{il}&=\partial_{j}g_{il}-g_{pl}\bar{\Gamma}^{p}_{ji}-g_{ip}\bar{\Gamma}^{p}_{jl}\\ -\bar{\nabla}_{l}g_{ij}&=-\partial_{l}g_{ij}+g_{pj}\bar{\Gamma}^{p}_{il}+g_{ip}\bar{\Gamma}^{p}_{lj}.\end{split}$
Taking the sum of both sides, we get
@@
[…]
@@
$\begin{split}\bar{\nabla}_{k}W^{k}_{ij}&-\bar{\nabla}_{i}W^{k}_{kj}+W^{p}_{ij}W^{k}_{kp}-W^{p}_{kj}W^{k}_{ip}\\ &=\partial_{k}\Gamma^{k}_{ij}-\partial_{i}\Gamma^{k}_{kj}+\Gamma^{p}_{ij}\Gamma^{k}_{kp}-\Gamma^{p}_{kj}\Gamma^{k}_{ip}\\ &~~~~~-(\partial_{k}\bar{\Gamma}^{k}_{ij}-\partial_{i}\bar{\Gamma}^{k}_{kj}+\bar{\Gamma}^{p}_{ij}\bar{\Gamma}^{k}_{kp}-\bar{\Gamma}^{p}_{kj}\bar{\Gamma}^{k}_{ip})\\ &=R_{ij}-\bar{R}_{ij}.\end{split}$
@@
Since $\bar{\nabla}\bar{g}=0$, we have
@@
where the term $g^{-1}*g^{-1}*g^{-1}*\bar{\nabla}h*\bar{\nabla}h$ is a contraction of three copies of $g^{-1}$ (i.e., $g$ with raised indices) and two copies of $\bar{\nabla}h=\bar{\nabla}g$. And, the term $(\bar{g}+h)^{-1}h\bar{g}^{-1}h\bar{g}^{-1}*\mathrm{Ric}_{\bar{g}}$ is the trace of $\mathrm{Ric}_{\bar{g}}$ with respect to $\left((\bar{g}+h)^{-1}h\bar{g}^{-1}h\bar{g}^{-1}\right)^{-1}$. Note that $\bar{g}+h$ is positive definite if $\|h\|_{\bar{g}}$ is small enough. Indeed, this formula follows by taking both sides of (26) in Proposition 7.2 with respect to $g=\bar{g}+h$ and using